Find the no. Of solution of equation   x2=2x.
Please reply and send solutions to mobie number   +919587748173.
Thank u.

Grade:12th Pass

3 Answers

palak singhal
9 years ago
Arun Kumar IIT Delhi
9 years ago
Hi

Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty
mycroft holmes
9 years ago
If x>0, then consider the function, f(x) = x1/x. This is monotonic in the intervals (0, e) and (e, infty)
 
Thus only two solutions exist to x1/x =2 1/2 i.e. x =2 and x=4
 
If x<0, then f(x) = 2x – xmonotonically increasing and continuous. Also, f(0)>0 and f(-1)<0 and so one more solution may be found in this interval. 
 
Thus three distinct solutions exist